demonstrate mathematically the rule of atomistic competition for energy commodity that each supplier determines its production according to the marginal cost equals price .



Afterwards,


Explain the demand side of the energy market situation

1
Expert's answer
2022-03-03T07:25:14-0500

Like in other markets, prices coordinate individual decisions on the supply and the demand side. At first sight, the model of an ideal market seems to apply to many energy markets: They can be clearly defined, products traded on them are highly homogeneous at least from a physical point of view, and many prices are transparent. If the number of independent suppliers is large, the corresponding energy market fits the model of perfect atomistic competition. This means that individual suppliers can only choose the quantity of energy Q they would like to offer (acting as so-called price takers). Let them maximize their per-period profit, i.e. the difference between revenue P.Q and total cost C(Q),

"\Pi(Q)= P.Q- C(Q)"

The solution to this problem can be found by setting the derivative of the profit function (1.1) with respect to the produced quantity Q equal to zero.

"\frac{\delta \Pi}{\delta Q}=\frac{\delta( P.Q)}{\delta Q}-\frac{\delta C}{\delta Q}=p^{'}-\frac{\delta C}{\delta Q}=0"

"\to C^{'}=\frac {\delta C}{\delta Q}=P{'}"

On the demand side, marginal willingness to pay derives from marginal utility of consumption. Demand for a good is triggered as long as its marginal utility exceeds the marginal cost of consumption (the market price in this simple model). In the case of energy, this is a derived demand because utility does not emanate directly from the consumption of energy but rather from the services associated with it, such as lighting, heating, use of appliances, and transportation. Therefore, the contribution of energy to the production of these services (its marginal productivity to be precise) has to be taken into account to determine the marginal utility of energy.